Three years from now, an uncle will be thrice as old as his nephew will be. Two years ago, he was four times as old as his nephew was then, How old are they now?

Answer by harpazo(655)

Let U = Uncle
Let n = Nephew
Here is your system of equations:
u + 3 = 3(n + 3)
u - 2 = 4(n - 2)
After solving the system by substitution, I got n = 12.
Let's plug n = 12 into either equation above to find u.
u + 3 = 3(12 + 3)
u + 3 = 36 + 9
u + 3 = 45
u = 45 - 3
u = 42
The nephew is 12 and the uncle is 42.
NOTE: 42 is thrice as old
42 + 3 = 3(12 + 3)
45 = 36 + 9
45 = 45
MORE PROVE:
u - 2 = 4(n - 2)
42 - 2 = 4(12 - 2)
40 = 4(10)
40 = 40
The nephew is 12 now and the uncle is 42 now.